# Barycentric Neural Networks and Length-Weighted Persistent Entropy Loss: A Green Geometric and Topological Framework for Function Approximation

### v2

@@ Abstract
While it is well-established that artificial neural networks are universal approximators for continuous functions on compact domains, many modern approaches rely on deep or overparameterized architectures that incur high computational costs. In this paper, a new type of small shallow neural network, called the Barycentric Neural Network ($\operatorname{BNN}$), is proposed, which leverages a fixed set of base points and their barycentric coordinates to define both its structure and its parameters. We demonstrate that our $\operatorname{BNN}$ enables the exact representation of continuous piecewise linear functions ($\operatorname{CPLF}$s), ensuring strict continuity across segments. Since any continuous function over a compact domain can be approximated arbitrarily well by $\operatorname{CPLF}$s, the $\operatorname{BNN}$ naturally emerges as a flexible and interpretable tool for function approximation. Beyond the use of this representation, the main contribution of the paper is the introduction of a new variant of persistent entropy, a topological feature that is stable and scale invariant, called the length-weighted persistent entropy ($\operatorname{LWPE}$), which is weighted by the lifetime of topological features. Our framework, which combines the $\operatorname{BNN}$ with a loss function based on our $\operatorname{LWPE}$, aims to provide flexible and geometrically interpretable approximations of nonlinear continuous functions in resource-constrained settings, such as those with limited base points for $\operatorname{BNN}$ design and few training epochs. Instead of optimizing internal weights, our approach directly optimizes the base points that define the $\operatorname{BNN}$. Experimental results show that our approach achieves superior and faster approximation performance compared to classical loss functions such as MSE, RMSE, MAE, and log-cosh. This efficiency aligns with the principles of Green AI, offering an effective trade-off between approximation performance and resource consumption. By unifying tools from geometry and topology within the neural network paradigm, our work provides a mathematically grounded yet computationally sustainable framework for function approximation.
@@ 1 Introduction
Function representation and approximation are fundamental problems in Machine Learning and scientific computing, underpinning a wide range of applications such as physical simulation, data modeling, and predictive analysis. A foundational result in this area is the Universal Approximation Theorem, which established that artificial neural networks (ANNs) are universal approximators, capable of approximating any continuous function over compact domains [15]. The theoretical study of ANNs as universal approximators has since evolved significantly, with further work demonstrating their capacity to approximate continuous and smooth functions effectively [10, 31, 16, 27]. More recent research has explored optimal approximation properties of sparsely connected deep networks [2] and the expressive power of ReLU-based architectures for function approximation [18, 23]. Although deep neural networks (DNNs) are powerful and are indeed known to be universal approximators, they typically require highly overparameterized architectures, large datasets, and significant computational resources for training. This often results in architectures (models) that are opaque, costly and energy-intensive. Such limitations are particularly problematic in settings with limited resources, such as embedded systems and edge computing, where lightweight and interpretable models are preferable.
@@
The growing tension between performance and interpretability has motivated the development of more structured, explainable, and computationally efficient models. This aligns with the emerging paradigm of Green Artificial Intelligence (Green AI) [3], which advocates for energy-efficient and sustainable machine learning systems. In this context, geometric and topological tools have gained increasing attention for their ability to structure information and capture global patterns that go beyond local pointwise accuracy. On the geometric side, barycentric coordinates offer a classical yet underutilized approach for function approximation. These coordinates allow points to be expressed as convex combinations of the vertices of a simplex, enabling a natural mechanism for piecewise linear interpolation. Historically applied in computer graphics and geometric design [30], barycentric interpolation has recently been extended to function approximation [22]. While [22] employs barycentric subdivisions, which lead to an exponential growth in the size of the triangulations, our approach does not rely on such subdivisions. These approaches demonstrate that fixed geometric structures can yield accurate and efficient representations, especially for continuous piecewise linear functions. Nested barycentric coordinate systems (NBCS) have been proposed to represent complex function spaces in a composable and interpretable manner [12]. Similarly, variational barycentric interpolation has been explored for mesh-based learning and shape optimization [7]. Despite their potential, geometric techniques, like barycentric coordinates, are rarely integrated into trainable models for function approximation, despite remaining decoupled from neural network architectures.
@@
On the topological side, topological methods have also been integrated into analysis and machine learning. Persistent homology, a core method in topological data analysis (TDA), and its associated summaries, like persistent entropy, have proven valuable for extracting global, deformation-invariant features from data. Initially used for feature extraction, TDA provides stable and noise-robust descriptors based on the birth and death of topological features such as connected components, loops, and voids, that can be used as input data for machine learning models across diverse tasks, such as characterizing idiotypic immune networks [25], analyzing similarities in piecewise linear functions and time series [26], and defining safety regions for robotic navigation [29]. More recently, the field of Topological Deep Learning (TDL) has emerged, incorporating persistent homology into the training of neural networks through differentiable topological layers and loss functions [33]. For instance, persistent homology and associated descriptors have been used to inject shape-aware constraints into neural networks, yielding improvements in tasks like anatomical segmentation [19], point cloud reconstruction [14], shape classification [24], and time series forecasting [17]. Recent advances in differentiable topological optimization [4, 32] have further improved the trainability and robustness of topological loss functions, making them viable components for neural network training pipelines. These topological losses aim to preserve global structural properties in learned functions, which classical loss functions like MSE tend to overlook. However, challenges persist, including gradient sparsity and computational inefficiency, which have limited their use in practice.
In this paper, we propose a novel, interpretable, and sustainable framework that combines geometric and topological methods for function approximation. We first introduce a new type of shallow neural network, the so-called Barycentric Neural Network ($\operatorname{BNN}$). Such neural network architecture enables efficient piecewise linear interpolation, making it well-suited for representing Continuous Piecewise Linear Functions ($\operatorname{CPLF}$s), ensuring strict continuity across segments. Since any continuous function over a compact domain can be approximated arbitrarily well by $\operatorname{CPLF}$s, the $\operatorname{BNN}$ naturally emerges as a flexible, interpretable, and resource-efficient tool for function approximation. We use barycentric coordinates to define the internal parameters of the proposed neural network architecture. This way, the $\operatorname{BNN}$ is fully determined by a set of trainable base points and their corresponding barycentric coordinates. These base points serve as the vertices of the simplices over which barycentric interpolation is performed. As a result, unlike traditional neural networks that require tuning internal weights through backpropagation, training a BNN reduces to optimizing the base points while the network’s parameters remain fixed by construction and depend exclusively on these base points. The number of base points can be specified in advance by the user and directly determines both the size and the expressiveness of the architecture: the larger the set, the more complex the architecture becomes and the greater its approximation capacity.
Our goal is to approximate a target continuous function, assuming we have limited resources, as only a small set of such base points and a few training epochs, based on the intuition that the more representative these points are, the better the approximation. To further improve approximation quality, we introduce a novel loss function based on a new topological summary statistic that we define, called length-weighted persistent entropy ($\operatorname{LWPE}$). This statistic is a modified version of classical persistent entropy [5], a stable and scale-invariant topological descriptor, enriched by weighting the contribution of each topological feature by its lifespan (i.e., persistence). We then employ this $\operatorname{LWPE}$-based loss function in our framework, guiding the optimization of the $\operatorname{BNN}$’s base points through topological information. Unlike classical loss functions such as MSE, RMSE, MAE, or log-cosh, this approach preserves not only pointwise accuracy but also the global topological structure of the target function. By combining the $\operatorname{BNN}$ with the $\operatorname{LWPE}$-based loss, our framework achieves superior function approximation compared to traditional loss functions when a limited number of base points and training epochs are used, aligning our methodology with the goals of Green AI. This is particularly useful in applications where topology carries semantic meaning, such as preserving peaks and valleys in economic indicators, motifs in biological signals, or capturing global trends in physical models.
@@
  • Barycentric Neural Network: We propose a scalable and shallow neural network, the so-called Barycentric Neural Network, for representing continuous piecewise linear functions. Its structure and parameters are fixed, based on a predefined set of points, called base points, and their barycentric coordinates.
  • Optimization on the base points: Instead of relying on backpropagation for weight tuning, the $\operatorname{BNN}$’s training is entirely based on optimizing the positions of its base points, yielding a geometrically and more efficient training process.
  • Topological loss function: We propose a novel topological summary statistic, the length-weighted persistent entropy ($\operatorname{LWPE}$), which serves as the basis for a new loss function. This formulation provides a topologically informed mechanism that enhances both convergence and the quality of function approximation using BNNs with a limited number of base points.
  • Experimental validation: We empirically validate the effectiveness of our approach ($\operatorname{BNN}$ trained with our $\operatorname{LWPE}$-based loss function) for function approximation using both synthetic and real-world data, showing superior performance and efficiency compared to using classical loss functions (e.g., MSE, RMSE, MAE and log-cosh), under resource-constrained settings, such as limited base points for $\operatorname{BNN}$ design and few training epochs.
By integrating geometric and topological tools into neural network design, this work bridges theoretical advances with practical utility, contributing to the ongoing development of efficient, interpretable and sustainable Machine Learning methodologies, while also enriching the emerging field of geometric and topological deep learning. In particular, it offers a practical alternative to traditional overparameterized deep networks in scenarios that demand structural guarantees and computational efficiency, such as function approximation under limited resources.
The remainder of this paper is organized as follows: Section 2 introduces the preliminary notions related to the different topics involved in the paper, such as barycentric coordinates, persistent homology and persistent entropy. Section 3 introduces our proposed variant of persistent entropy, the length-weighted persistent entropy. Section 4 details how barycentric neural networks are defined. In Section 5, we present the experiments conducted, including an explanation of how our length-weighted persistent entropy-based loss function is defined, and how we use it for function approximation in combination with the barycentric neural network, along with the results obtained. Finally, conclusions, limitations and future work are given in Section 6.
@@ 2 Preliminaries
This section introduces the main concepts and establishes some notations that are used throughout the paper. We begin by introducing $\operatorname{CPLF}$, followed by the key concepts of computational topology employed in this work, including simplicial complexes, barycentric coordinates, persistent homology, and persistent entropy. A general introduction to computational topology can be found, for example, in [8, 21].
@@
Fixed $d\in\mathbb{N}$, a Continuous Piecewise Linear Function ($\operatorname{CPLF}$) is a function $h:C\subseteq{\mathbb{R}}^{d}\to{\mathbb{R}}$, where the domain $C$ is the union of specific regions $|\sigma_{0}|,\dots,|\sigma_{n}|$ (called $d$-simplices, defined later). On each simplex $|\sigma_{k}|$, the function $h$ is given by a linear expression: $h(p)=m_{k}\cdot p+c_{k}$, for some $m_{k}\in{\mathbb{R}}^{d}$. $c_{k}\in{\mathbb{R}}$, and $p\in|\sigma_{k}|$. Importantly, the function $h$ is required to be globally continuous, ensuring that the linear expressions agree on overlapping regions, that is, $m_{i}\cdot p+c_{i}=m_{j}\cdot p+c_{j}$ whenever $p\in|\sigma_{i}|\cap|\sigma_{j}|$.
@@
In this paper, only $\operatorname{CPLF}$ functions $h:[A,B]\subset\mathbb{R}\to\mathbb{R}$, where $[A,B]$ is the domain interval with endpoints $A,B\in\mathbb{R}$, will be considered for experiments in Section 5. In that case, for some $n\in\mathbb{N}$, the domain interval $[A,B]$ of $h$ is divided into $n+1$ subintervals $a_{i},b_{i}]$, where $A=a_{0}<b_{0}=a_{1}<\cdots<b_{n-1}=a_{n}<b_{n}=B$. The function $h$ is continuous at the joining points, that is, $m_{i-1}\cdot b_{i-1}+c_{i-1}=m_{i}\cdot a_{i}+c_{i}$. Any given continuous function $f:[A,B]\subset\mathbb{R}\to\mathbb{R}$ can be approximated as closely as desired by a $\operatorname{CPLF}$ $h:[A,B]\to\mathbb{R}$ by partitioning $[A,B]$ into a a sufficiently large finite number of segments.
@@
Any compact subset $\mathcal{C}$ of $\mathbb{R}^{d}$ can be approximated by a simplicial complex (see, for example, [20, Chapter 1]). The combinatorial structure of such complexes allows for efficient computation of topological information using discrete algorithms. Given $d\in\mathbb{N}$, a (finite, $d$-dimensional, embedded in $\mathbb{R}^{d}$) simplicial complex $\mathcal{K}$ consists of:
@@
Barycentric coordinates naturally connect geometry to topology through simplicial complexes. In any dimension $d$, barycentric coordinates $(t_{0},t_{1},\ldots,t_{d})$ describe the location (position) of a point $p\in\mathbb{R}^{d}$ within the vertices of a $d$-simplex. Formally, for any simplex $\sigma=\{v_{0},v_{1},\ldots,v_{d}\}$ with vertices $v_{i}\in\mathbb{R}^{d}$ in general position and any point $p\in|\sigma|\subset\mathbb{R}^{d}$, there exists unique real values $(t_{0},t_{1},\ldots,t_{d})$ such that:
@@
This representation allows a continuous interpolation across the simplices of a simplicial complex.
The barycentric coordinates of a point $p=(p_{1},\dots,p_{d})^{\scriptscriptstyle T}\in\mathbb{R}^{d}$ with respect to a simplex $\sigma=\{v_{0},v_{1},\dots$, $v_{d}\}$ can always be computed. If we write the coordinates of the point $p$ and the vertices $v_{i}$ in column form, then we have:
$(\begin{array}[]{cccc}t_{0}&t_{1}&\cdots&t_{d}\end{array})^{\scriptscriptstyle T}=\left(\begin{array}[]{cccc}v_{0}&v_{1}&\cdots&v_{d}\\ 1&1&\cdots&1\end{array}\right)^{-1}\cdot\left(\begin{array}[]{c}p\\ 1\end{array}\right)\,.$
This formulation ensures that $\sum_{i=0}^{d}t_{i}=1$ and that $t_{i}\geq 0$ for all $i\in\{0,1,\dots,d\}$ if and only if $p$ lies in $|\sigma|$. Besides, the inverse of the $(d+1)\times(d+1)$-matrix $\left(\begin{array}[]{cccc}v_{0}&v_{1}&\cdots&v_{d}\\ 1&1&\cdots&1\end{array}\right)$ always exists because $\{v_{0},v_{1},\dots,v_{d}\}$ are in general position.
@@
The barycentric coordinates of a point $x\in\mathbb{R}$ with respect to an interval (i.e., a 1-simplex) $a,b]\subset\mathbb{R}$ are (t,1t)(t,1-t) with $t=\frac{x-a}{b-a}$. Then, $(1-t,t)\in[0,1]^{2}$ if and only if $x\in[a,b]$.
Lower Star Filtration
Fixed $d\in\mathbb{N}$, let us consider a simplicial complex $\mathcal{K}$ with real values assigned to its vertices $V\subset\mathbb{R}^{d}$, denoted by $g\colon\mathbb{R}^{d}\to\mathbb{R}$. The lower star of a vertex $u\in V$ is the subcomplex that consists of all the simplices $\sigma$ of $\mathcal{K}$ such that $u$ is the minimum vertex of $\sigma$, that is $u\in\sigma$ and $g(u)\leq g(v)$ for any $v\in\sigma$. Assuming $V$ is composed by vertices $\{u_{0},u_{1},\dots,u_{m}\}$ with distinct values $g(u_{0})<g(u_{1})<\ldots<g(u_{m})$, the lower star filtration of $\mathcal{K}$ is the increasing sequence of simplicial complexes $\mathcal{K}(1)\subset\mathcal{K}(2)\subset\cdots\subset\mathcal{K}(n)=\mathcal{K}$ where $\mathcal{K}(i)$ is the lower star of vertex $u_{i}$. See [8, Chapter VI]. As we will see later, this filtration will be used to compute homology, and consequently persistent entropy, as well as our topological loss function, which guides the optimization of the base points in $\operatorname{BNN}$s.
@@
Persistent homology is a tool for studying data geometry and connectivity across scales. Using this tool and given a filtered simplicial complex (e.g., lower star filtration), one can effectively compute $q$-dimensional topological features at different scales, tracking features as they appear (birth) and disappear (death). Persistence diagrams and barcodes are visual representations used in the study of persistent homology and represent the birth and death (end) of topological features across scales. A $q$-dimensional persistence diagram plots points $(b,e)$ corresponding to the lifespans of $q$-dimensional topological features, while a barcode represents them as a bar starting at $b$ and ending at $e$. Specifically, a $q$-dimensional persistence diagram $D=\{(b_{i},e_{i})\mid i\in I\}$ satisfies that $b_{i}\leq e_{i}$ for all $i\in I$, and $I$ is the index set that identifies the pairs $(b_{i},e_{i})$ in $D$. Each pair $(b_{i},e_{i})$ corresponds to a $q$-dimensional topological feature that appears at time $b_{i}$ (birth) and disappears at time $e_{i}$ (death, i.e., end) as the filtration progresses. Both persistence diagrams and barcodes facilitate the understanding of feature persistence in data. Fig. 1 shows a persistence diagram (center) alongside its corresponding barcode (right), illustrating the persistence of topological features in the data.
The number of bars of the barcode associated with the lower star filtration of a simplicial complex $\mathcal{K}$ is less than or equal to half of the number of vertices of the simplicial complex $\mathcal{K}$ plus 1. Given the lower star filtration on a simplicial complex $\mathcal{K}$ with vertex set $V$, local critical points are always located at the vertices of $V$ and can be effectively computed from the lower link of each vertex $v$ of $V$, which consists of the simplices of $\mathcal{K}$ in the closed lower star that do not belong to the lower star of $v$. We call $v$ a local critical vertex of index $q$ if its lower link has the reduced homology of the (q1)(q-1)-sphere. Then, $q$-dimensional persistence diagrams pair critical vertices of index $q$ (when topological features are born) with critical vertices of index $q+1$ (when topological features die). A detailed description can be found in [9, Chapter VI].
@@
Persistent entropy reflects the uniformity of bar lengths in a persistence barcode. Its maximum is attained when all bars have equal length, in which case it is given by $\operatorname{PE}=\ln(\#I)$, where #I\#I denotes the cardinality of $I$ (i.e., the number of bars in the persistence diagram $D$). Conversely, its minimum value is $\operatorname{PE}=0$, which occurs when the barcode contains a single long bar and the remaining bars have negligible length. Moreover, the greater the disparity among bar lengths, the closer the value of persistent entropy is to its minimum.
The stability results on persistence diagrams [6] and persistent entropy [1, 26] provide a robust foundation for quantifying the stability of topological features under small perturbations. Specifically, it is ensured that small changes in the input correspond to bars of short lengths in the persistence barcode and small, bounded changes in the persistent entropy. This property is particularly valuable in applications where small variations in data or model outputs, such as those encountered in numerical approximations or machine precision errors, might otherwise introduce significant uncertainties in analytical results. As an illustration, the persistent entropy computed in the example shown in Fig. 1 is 1.3806.

### v3

@@ Abstract
While artificial neural networks are known as universal approximators for continuous functions, many modern approaches rely on overparameterized architectures with high computational cost. In this work, we introduce the Barycentric Neural Network ($\operatorname{BNN}$): a compact shallow architecture that encodes both structure and parameters through a fixed set of base points and their associated barycentric coordinates. We show that the $\operatorname{BNN}$ enables the exact representation of continuous piecewise linear functions ($\operatorname{CPLF}$s), ensuring strict continuity across segments. Given that any continuous function on a compact domain can be uniformly approximated by $\operatorname{CPLF}$s, the $\operatorname{BNN}$ emerges as a flexible and interpretable tool for function approximation.
To enhance geometric fidelity in low-resource scenarios, such as those with few base points to create $\operatorname{BNN}$s or limited training epochs, we propose length-weighted persistent entropy ($\operatorname{LWPE}$): a stable variant of persistent entropy. Our approach integrates the $\operatorname{BNN}$ with a loss function based on $\operatorname{LWPE}$ to optimize the base points that define the $\operatorname{BNN}$, rather than its internal parameters. Experimental results show that our approach achieves superior and faster approximation performance compared to standard losses (MSE, RMSE, MAE and LogCosh), offering a computationally sustainable alternative for function approximation.
@@ 1 Introduction
Function representation and approximation are fundamental problems in machine learning and scientific computing, underpinning a wide range of applications such as physical simulation, data modeling, and predictive analysis. A cornerstone of this field is the Universal Approximation Theorem, which guarantees that artificial neural networks (ANNs) can approximate any continuous function on a compact domain [15]. Since then, theoretical advances have deepened our understanding of ANN’s approximation capabilities, particularly for smooth functions [10, 31, 16], and have highlighted the expressive power of ReLU-based architectures for representing piecewise linear structures [18, 24]. More recent work has also explored the approximation efficiency [2] and the role of depth versus width in function approximation [28]. However, modern deep neural networks (DNNs) typically require highly overparameterized architectures, large datasets, and extensive computational resources. This leads to architectures (models) that are not only energy-intensive and opaque, but also poorly suited for resource-constrained environments, such as embedded systems or edge devices, where lightweight and efficient solutions are essential.
@@
This tension between performance and sustainability has spurred interest in methods that can achieve high performance with reduced computational overhead. In particular, the emerging paradigm of Green AI [3] calls for machine learning methods that balance performance with computational and environmental cost. Within this context, geometric tools like barycentric coordinates offer a promising yet underexplored avenue. These coordinates allow points to be expressed as convex combinations of the vertices of a simplex, enabling a natural mechanism for piecewise linear interpolation. Historically applied in computer graphics and geometric design [30], barycentric interpolation has recently been extended to function approximation, providing interpretability and expressivity [22, 23]. However, these methods still rely on barycentric subdivisions, which could limit flexibility. In contrast, our approach does not rely on subdivisions. In addition, nested barycentric coordinate systems (NBCS) have been proposed to represent complex function spaces in an interpretable manner [12]. Similarly, variational barycentric interpolation has been explored for mesh-based learning and shape optimization [7].
@@
Besides, topological methods have also been integrated into data analysis and machine learning. Persistent homology, a core method in topological data analysis (TDA), and its associated summaries, like persistent entropy, have proven valuable for extracting global, deformation-invariant features from data. Initially used for feature extraction, TDA provides stable and noise-robust descriptors based on the birth and death of topological features (e.g., connected components, loops) across scales, that can be used as input data for machine learning models across diverse tasks, such as immune networks characterization [26] and time series comparison [27]. More recently, the field of topological deep learning (TDL) has emerged, incorporating persistent homology into the training of neural networks via differentiable topological layers and loss functions [32]. For instance, persistent homology and associated descriptors have been used to inject shape-aware constraints into neural networks, yielding improvements in tasks like anatomical image segmentation [19], point cloud reconstruction [14], shape classification [25], and time series forecasting [17]. Recent advances in differentiable topological optimization [4] have further enhanced the trainability and robustness of topological loss functions, making them viable components for neural network training pipelines. However, challenges such as gradient sparsity and computational overhead continue to hinder widespread adoption.
In this paper, we bridge geometry and topology to propose a novel, interpretable, and sustainable framework for function approximation under limited resources. Our approach centers on the Barycentric Neural Network ($\operatorname{BNN}$), a shallow, compact architecture that exactly represents continuous piecewise linear functions ($\operatorname{CPLF}$s) with guaranteed continuity across regions. Since any continuous function over a compact domain can be uniformly approximated by $\operatorname{CPLF}$s, the $\operatorname{BNN}$ naturally emerges as a flexible, interpretable, and efficient tool for function approximation. Crucially, the $\operatorname{BNN}$’s structure and parameters are fully determined by a set of trainable base points and their barycentric coordinates. Rather than optimizing internal weights, training reduces to optimizing the positions of these base points, turning the learning problem into a geometric one. The number of base points directly controls model capacity: fewer points yield lightweight models, while more enable higher fidelity, offering a natural trade-off between expressiveness and resource usage. To guide this geometric optimization, we introduce length-weighted persistent entropy ($\operatorname{LWPE}$), a new topological summary that extends classical persistent entropy by weighting each feature’s contribution by its persistence (lifetime). By emphasizing features that persist across multiple scales, this approach naturally prioritizes the most structurally important characteristics of the target function. This yields a stable, scale-invariant descriptor. We incorporate $\operatorname{LWPE}$ into a topologically informed loss function that preserves not only pointwise accuracy but also the global shape of the target function, something traditional losses like $\operatorname{MSE}$, $\operatorname{RMSE}$, $\operatorname{MAE}$, or LogCosh often neglect.
Our framework, $\operatorname{BNN}$ trained with $\operatorname{LWPE}$-based loss, is particularly effective in low-resource settings, such as when only a small number of base points or few training epochs are available. Experimental results on synthetic and real-world data demonstrate that our method achieves faster convergence and superior approximation quality compared to standard losses when limited resources are available, aligning it with the goals and principles of Green AI. This is especially valuable in domains where topological fidelity matters, such as preserving peaks and valleys in economic time series, motifs in biological signals, or global trends in physical simulations.
@@
  • Barycentric Neural Network: A shallow and scalable architecture that exactly represents $\operatorname{CPLF}$s. Its structure and parameters are fixed, based on a trainable and predefined set of points and their barycentric coordinates.
  • Geometric training via base points optimization: The $\operatorname{BNN}$’s training is entirely based on optimizing the positions of its base points, yielding a geometrically and more efficient training process.
  • Lenght-weighted persistent entropy: A stable topological summary that weights features by persistence, enabling a more informative topological loss.
  • Experimental validation under resource constraints: Comprehensive experiments showing that $\operatorname{BNN}$s trained with $\operatorname{LWPE}$ outperform classical losses (such as $\operatorname{MSE}$, $\operatorname{RMSE}$, $\operatorname{MAE}$ or LogCosh) in function approximation performance and convergence speed when base points and training epochs are limited.
By unifying geometric representation with topological regularization, our work contributes to the growing fields of topological deep learning and Green AI, while offering a practical, sustainable, and interpretable alternative to overparameterized models in function approximation tasks.
The remainder of this paper is organized as follows: Section 2 reviews key concepts involved in the paper, such as barycentric coordinates, persistent homology and persistent entropy. Section 3 introduces length-weighted persistent entropy. Section 4 details how barycentric neural networks are defined. In Section 5, we present the experiments conducted, including an explanation of how our length-weighted persistent entropy-based loss function is defined, and how we use it for function approximation in combination with the barycentric neural network in low-resource scenarios, along with the results obtained. Finally, conclusions and future work are given in Section 6.
@@ 2 Preliminaries
This section introduces the main concepts and establishes some notations that are used throughout the paper. We begin by introducing $\operatorname{CPLF}$, followed by the key concepts of computational topology employed in this work, including simplicial complexes, barycentric coordinates, persistent homology, and persistent entropy. A general introduction to computational topology can be found, we refer the reader to [8, 21].
@@
Given $d\in\mathbb{N}$, a continuous piecewise linear function ($\operatorname{CPLF}$) is a function $h:C\subseteq{\mathbb{R}}^{d}\to{\mathbb{R}}$, where the domain $C$ is the union of specific regions $|\sigma_{0}|,\dots,|\sigma_{n}|$ (called $d$-simplices, defined later). On each simplex $|\sigma_{k}|$, the function is affine:
$h(p)=m_{k}\cdot p+c_{k},\quad\text{for }p\in|\sigma_{k}|,$
with $m_{k}\in\mathbb{R}^{d}$ and $c_{k}\in\mathbb{R}^{d}$. Global continuity is enforced by requiring that adjacent linear pieces agree on shared faces: if $p\in|\sigma_{i}|\cap|\sigma_{j}|$, then $m_{i}\cdot p+c_{i}=m_{j}\cdot p+c_{j}$.
@@
In our experiments (Section 5), we restrict attention to one-dimensional domains: $h\colon[A,B]\subset\mathbb{R}\to\mathbb{R}$. Here, the domain is partitioned into subintervals $a_{i},b_{i}]$ with $A=a_{0}<b_{0}=a_{1}<\cdots<b_{n-1}=a_{n}<b_{n}=B$, and continuity at breakpoints ensures $m_{i-1}\cdot b_{i-1}+c_{i-1}=m_{i}\cdot a_{i}+c_{i}$. Any given continuous function $f:[A,B]\subset\mathbb{R}\to\mathbb{R}$ can be approximated as closely as desired by a $\operatorname{CPLF}$ $h:[A,B]\to\mathbb{R}$ by partitioning $[A,B]$ into a a sufficiently large finite number of segments.
@@
Given $d\in\mathbb{N}$, a (finite, $d$-dimensional, embedded in $\mathbb{R}^{d}$) simplicial complex $\mathcal{K}$ consists of:
@@
Any compact subset $\mathcal{C}$ of $\mathbb{R}^{d}$ can be approximated by a simplicial complex (see, for example, [20, Chapter 1]). The combinatorial structure of such complexes allows for efficient computation of topological information using discrete algorithms.
@@
Barycentric coordinates naturally connect geometry to topology through simplicial complexes. In any dimension $d$, barycentric coordinates $(t_{0},t_{1},\ldots,t_{d})$ describe the location (position) of a point $p\in\mathbb{R}^{d}$ within the vertices of a $d$-simplex, ensuring that $t_{i}\geq 0$ for all $i\in\{0,1,\dots,d\}$ if and only if $p$ lies in $|\sigma|$. Besides, barycentric coordinates representation allows a continuous interpolation across the simplices of a simplicial complex. Formally, for any simplex $\sigma=\{v_{0},v_{1},\ldots,v_{d}\}$ with vertices $v_{i}\in\mathbb{R}^{d}$ in general position and any point $p\in|\sigma|\subset\mathbb{R}^{d}$, there exists unique real values $(t_{0},t_{1},\ldots,t_{d})$ such that:
@@
This representation allows a continuous interpolation across the simplices of a simplicial complex. Furthermore, unlike in [22], where the barycentric coordinates $(t_{0},t_{1},\dots,t_{d})$ of a point $p\in\mathbb{R}^{d}$ with respect to a simplex $\sigma=\{v_{0},v_{1},\dots$, $v_{d}\}$ are computed using a matrix inverse, we instead employ the following equivalent formulation, which expresses the coordinates as determinants via Cramer’s rule:
$t_{i}=\frac{\lambda_{i}}{\sum_{j=0}^{d}\lambda_{j}},\qquad\lambda_{i}=\det\left(\begin{array}[]{ccccccc}v_{0}&\cdots&v_{i-1}&p&v_{i+1}&\cdots&v_{d}\\ 1&\cdots&1&1&1&\cdots&1\end{array}\right),\qquad i=0,1,\dots,d.$
This formulation gives explicit, differentiable, local, and interpretable expressions for barycentric coordinates. Moreover, it can be implemented as a neural network layer where each $\lambda_{i}$ serves as a linear score, all scores are computed independently, and the final normalization acts as a softmax operator without the exponential.
@@
The barycentric coordinates of a point $x\in\mathbb{R}$ with respect to an interval (i.e., a 1-simplex) $a,b]\subset\mathbb{R}$ are (1t,t)(1-t,t) with $t=\frac{x-a}{b-a}$. Then, $(1-t,t)\in[0,1]^{2}$ if and only if $x\in[a,b]$.
Lower-Star Filtration
Fixed $d\in\mathbb{N}$, let us consider a simplicial complex $\mathcal{K}$ with real values assigned to its vertices $V\subset\mathbb{R}^{d}$, denoted by $g\colon\mathbb{R}^{d}\to\mathbb{R}$. Assuming $V$ is composed by vertices $\{u_{0},u_{1},\dots,u_{m}\}$ with distinct values $g(u_{0})<g(u_{1})<\ldots<g(u_{m})$, the lower-star of a vertex $u\in V$ is the subcomplex that consists of all the simplices $\sigma$ of $\mathcal{K}$ such that $u$ is the minimum vertex of $\sigma$, that is $u\in\sigma$ and $g(u)<g(v)$ for any $v\in\sigma$. The lower-star filtration of $\mathcal{K}$ is the increasing sequence of simplicial complexes $\mathcal{K}(1)\subset\mathcal{K}(2)\subset\cdots\subset\mathcal{K}(n)=\mathcal{K}$ where $\mathcal{K}(i)$ is the lower-star of vertex $u_{i}$. See [8, Chapter VI]. As we will see later, this filtration will be used to compute persistent homology, and consequently the topological loss based on $\operatorname{LWPE}$ used in the experiments of this paper.
@@
Persistent homology is a tool for studying data geometry and connectivity across scales. Using this tool and given a filtered simplicial complex (e.g., lower-star filtration), one can effectively compute $q$-dimensional topological features at different scales, tracking features as they appear (birth) and disappear (death). Persistence diagrams and barcodes are visual representations used in the study of persistent homology and represent the birth and death (end) of topological features across scales. A $q$-dimensional persistence diagram plots points $(b,e)$ corresponding to the lifespans of $q$-dimensional topological features, while a barcode represents them as a bar starting at $b$ and ending at $e$. Specifically, a $q$-dimensional persistence diagram $D=\{(b_{i},e_{i})\mid i\in I\}$ satisfies that $b_{i}\leq e_{i}$ for all $i\in I$, and $I$ is the index set that identifies the pairs $(b_{i},e_{i})$ in $D$. Each pair $(b_{i},e_{i})$ corresponds to a $q$-dimensional topological feature that appears at time $b_{i}$ (birth) and disappears at time $e_{i}$ (death, i.e., end) as the filtration progresses. Both persistence diagrams and barcodes facilitate the understanding of feature persistence in data. Fig. 1 shows a persistence diagram (center) alongside its corresponding barcode (right), illustrating the persistence of topological features in the data.
The number of bars of the barcode associated with the lower-star filtration of a simplicial complex $\mathcal{K}$ is less than or equal to half of the number of vertices of the simplicial complex $\mathcal{K}$ plus 1. Given the lower-star filtration on a simplicial complex $\mathcal{K}$ with vertex set $V$, local critical points are always located at the vertices of $V$ and can be effectively computed from the lower link of each vertex $v$ of $V$, which consists of the simplices of $\mathcal{K}$ in the closed lower-star that do not belong to the lower-star of $v$. We call $v$ a local critical vertex of index $q$ if its lower link has the reduced homology of the (q1)(q-1)-sphere. Then, $q$-dimensional persistence diagrams pair critical vertices of index $q$ (when topological features are born) with critical vertices of index $q+1$ (when topological features die). A detailed description can be found in [9, Chapter VI].
@@
Persistent entropy reflects the uniformity of bar lengths in a persistence barcode. Its maximum is attained when all bars have equal length, in which case it is given by $\operatorname{PE}=\ln(\#I)$, where #I\#I denotes the cardinality of $I$ (i.e., the number of points in the persistence diagram $D$). Conversely, its minimum value is $\operatorname{PE}=0$, which occurs when the barcode contains a single long bar and the remaining bars have negligible length. Moreover, the greater the disparity among bar lengths, the closer the value of persistent entropy is to its minimum.
The stability results on persistence diagrams [6] and persistent entropy [1, 27] provide a robust foundation for quantifying the stability of topological features under small perturbations. This property is particularly valuable in applications where small variations in data or model outputs, such as those encountered in numerical approximations or machine precision errors, might otherwise introduce significant uncertainties in analytical results. As an illustration, the persistent entropy computed in the example shown in Fig. 1 is 1.3806.